A 23 L container is filled with helium and the pressure is 154 atm and the temperature is 294 K. What volume can be filled in liters when the temperature is 18° C and the atmospheric pressure is 754 torr?
1
Expert's answer
2019-03-25T07:30:54-0400

"18C+273.15=291.15K"

"754torr*133.322=100525Pa"

"154atm*101325=1.56*10^{7}"

"n(He)=\dfrac{PV}{RT}=\dfrac{23L*1.56*10^{7}}{8.31J/mol*K}=146860.188mol"

"V=\dfrac{nRT}{P}=\dfrac{146860.188mol*8.31J/mol*K*291.15K}{100525Pa}=3534.661m^{3}=3534661L"